OpenStudy (babynini):

Prove the identity (tanx)/(1-cosx) = cscx(1+secx)

OpenStudy (babynini):

@ganeshie8 you free?

OpenStudy (babynini):

@Empty

OpenStudy (babynini):

@jim_thompson5910

OpenStudy (babynini):

@Nnesha

OpenStudy (babynini):

@freckles

OpenStudy (babynini):

I've gotten up to (sinx/cosx )(1/(1-cosx)

OpenStudy (babynini):

@IrishBoy123

rvc (rvc):

@Babynini u here?

rvc (rvc):

you can take the RHS to prove the LHS it seems easy from right to left

rvc (rvc):

change \(\Large\rm cscx\) to \[\Large\rm \frac{ 1 }{ sinx }\] and \(\Large\rm secx\) to \(\Large\rm\frac{1}{cos(x)}\)

OpenStudy (anonymous):

Let's solve for a. tanx1−cosx=cscx(1+s(2.718282)cx) Step 1: Multiply both sides by -cosx+1. antx=−2.718282c4os3x3−c3os2x2+2.718282c3s2x2+c2sx Step 2: Divide both sides by ntx. antxntx=−2.718282c4os3x3−c3os2x2+2.718282c3s2x2+c2sxntx a=−2.718282c4os3x2−c3os2x+2.718282c3s2x+c2snt Answer: a=−2.718282c4os3x2−c3os2x+2.718282c3s2x+c2snt

OpenStudy (babynini):

@rvc sorry, I wasn't online. So..what next?